OpenStudy (anonymous):

Find the surface area of a cylinder with diameter 22 in and height 15 in.

OpenStudy (jdoe0001):

\(\large { \textit{total surface area of a cylinder}=2\pi r(h+r) \\ \quad \\ r=radius\to \frac{diameter}{2}\qquad h=height }\)

OpenStudy (anonymous):

so 69 plus 15?

OpenStudy (wolf1728):

Surface Area Cylinder = (2 • π • r²) + (2 • π • r • height) radius =11 height = 15 Surface Area Cylinder = 2 * PI * 11² + (2*PI*11*15) Surface Area Cylinder = 760.2654221687 + 1,036.7255756846 Total = 1796.9909978534 or about 1,797 (rounded) and to check the calculations, a cylinder calculator: http://1728.org/diam.htm